OpenStudy (3psilon):

Determine where, if any where, the tangent line to f(x) =x^3 - 5x^2 +x is parallel to the line y = 4x +23. I derived it to be 3x^2 - 10x + 1 . But not sure what to do next

OpenStudy (anonymous):

well u should solve the equation \[3x^2 - 10x + 1=4\]why?

OpenStudy (3psilon):

Why not = to 4x + 23?

OpenStudy (anonymous):

actually derivative of a function in the given point equals to the slope of tangent line in that point

OpenStudy (3psilon):

Wait I don't quite understand

OpenStudy (anonymous):

what @mukushla means, is that the slope of the y = 4x +23 is equal to 4. The derivative you took is the slope of the tangent line. Si since you looking for a paralel to y = 4x +23, this derivative should be equal to 4 . So solve 3x^2 - 10x + 1 =4 for x to find which x gives you this slope.

OpenStudy (3psilon):

@myko I got X1 = -2.7 and x2=3.6

OpenStudy (3psilon):

Am I on the right track @mukushla

OpenStudy (anonymous):

x1 is not right i think

OpenStudy (anonymous):

so..at the points (x1,y1),(x2,y2)........ the curve has two tangents parallel to the given .line......... right??! @mukushla

OpenStudy (3psilon):

But I use the quadratic and keep getting that answer

OpenStudy (anonymous):

yes anusha

OpenStudy (anonymous):

u have\[3x^2-10x-3=0\]it gives: \[x_1=-0.27 \ \ \ x_2=3.61\]

OpenStudy (3psilon):

Oh sorry mine was a typo. I meant -.27 But what do you do next?

OpenStudy (3psilon):

@mukushla

OpenStudy (anonymous):

That's it. You found two points where tangent line to the curve f(x) =x^3 - 5x^2 +x has same slope (is parallel) to y = 4x +23